OpenStudy (tennistar):

What is the y-coordinate of the y-intercept of the line that passes through the points (-4, -4) and (4, 8)?

OpenStudy (jackellyn):

y=mx+b like in your previous question m=slope y= y intercept find the slope m= y2-y1/x2-x1 m=8-(-4)/4-(-4)=12/8=3/2 then substitute using this formula (y-y1)=m(x-x1) using either of the points y-(-4)=3/2(x-(-4)) y+4 = 3/2(x+4) y+4=3/2x + 6 y = 3/2x +6-4 y= 3/2x + 2 So the y intercept is (0,2)